Please need help. Thanks all the tutors here.
Determine the first 5 susesion terms of:
an=2n+5

Answer by stanbon(75887)

Determine the first 5 succesive terms of:
a(n) = 2n + 5
-------
a(1) = 2+5 = 7
a(2) = 2*2+5 = 9
a(3) = 2*3+5 = 11
a(4) = 2*4+5 = 13
a(5) = 2*5+5 = 15
----------------
Cheers,
Stan H.
---------------